### v1

# Space as Time Through Neuron Position Learning

Abstract
Biological neural networks exist in physical space where distance determines communication delays: a fundamental space-time coupling absent in most artificial neural networks. While recent work has separately explored spatial embeddings and learnable synaptic delays in spiking neural networks, we unify these approaches through a novel neuron position learning algorithm where delays relate to the Euclidean distances between neurons. We derive gradients with respect to neuron positions and demonstrate that this biologically-motivated constraint acts as an inductive bias: networks trained on temporal classification tasks spontaneously self-organize into local, small-world topologies with modular structure emerging under distance-dependent connection costs. Remarkably, we observe unprompted functional specialization aligned with spatial clustering without explictly enforcing it. These findings lay the groundwork for networks in which space and time are intrinsically coupled, offering new avenues for mechanistic interpretability, biologically inspired modelling, and efficient implementations.
1 Introduction
Brain networks face fundamental trade-offs between metabolic costs and information processing capabilities. Networks must minimise the costs of building and maintaining connections in physical space while optimising for computational capability. This constraint shapes brain organisation across species and may explain why brains converge on similar structural solutions (Achterberg et al., 2023), including sparse, small-world (Van den Heuvel et al., 2016) architectures and modular (Kaiser & Hilgetag, 2004) organisation.
To analyse the role of space, recent studies have started focusing on the spatial embedding of biological neural networks (Achterberg et al., 2023; Erb et al., 2025; Sheeran et al., 2024; Vasilache et al., 2025) – a critical dimension largely overlooked in computational models. However, these studies do not investigate the relationship between this spatial embedding and time – despite the spatial locations of neurons and their relative timing of information processing being naturally intertwined in the brain. Biological neural networks exist in physical space, where distance translates directly into time through conduction delays—a ‘space as time’ relationship (Voges & Perrinet, 2010; Izhikevich & Hoppensteadt, 2009) that is fundamental to brain organisation. The cost of connections is proportional to their length in three-dimensional space, and conduction delays vary dramatically across the nervous system due to differences in path length, diameter, and can even be optimised through myelin plasticity (Bonetto et al., 2021), i.e the brain could implement something akin to delay learning.
One of the first algorithms that employed a form of delay learning used multapses with various delays between neuron pairs and a corresponding trainable weight (Bohte et al., 2002). Since the delays could not be explicitly optimised, this algorithm could be understood more as a form of structure learning where delays are fixed and the optimal values are found through optimising the connectivity weights. The relationship between network structure and delays is still of interest, as Hammouamri et al. highlighted the benefits of delay learning in sparse networks, and Mészáros et al. (2024) observed that structural plasticity in networks with delays leads to better performance than explicit delay learning. Regardless, recent methods focus on gradient-based approaches (Hammouamri et al., ; Göltz et al., 2024; Sun et al., 2023), with Mészáros et al. (2024) introducing delay learning for recurrent connections.
The intertwinement of space and time creates natural constraints on both the structure and the dynamics of neural networks. Spatial clustering minimises costly long-range connections while creating fast communication within local regions and slower communication between distant areas. Thus, spatial organisation inherently generates temporal separation, where different brain regions operate on different timescales determined by their connectivity patterns, yielding hierarchically organised areas with increased temporal receptive windows (Chang et al., 2022). Evolution has exploited this relationship, as exemplified by the barn owl’s sound localisation system, which relies on precisely tuned delays arising from spatial circuit organisation (Carr & Konishi, 1988; Ghosh et al., 2025). These observations have not yet been modelled as, in most cases, neural networks discard spatial and temporal information. Even Spiking Neural Networks (SNNs) typically only adapt weights and biases – failing to exploit the spatial embedding and delay structure that biological systems have optimized over evolution.
Achterberg et al. (2023) proposed an approach for studying spatial structure in ANNs – giving us a way to connect them to temporal processing. However, their models only saw the cost of the connections, rather than the benefit of their structure. Erb et al. (2025) and Vasilache et al. (2025) explored position learning through distance-based synaptic weights, but disregarded the temporal aspects of space.
Now the foundations for both spatial aspects (through spatially embedding neurons) and the temporal aspects (through delay learning) have been laid, a natural next step is combining these into a single framework. Here, we do just that and derive a gradient-based learning algorithm for SNNs, where synaptic delays are determined by the Euclidean distance between two connected neurons. Furthermore, we derive the gradients of loss functions with respect to the positions of neurons, introducing a neural network shape learning algorithm. We study the various effects of the algorithm through training recurrent SNNs on the Spiking Heidelberg Digits (SHD) (Cramer et al., 2020) classification task. We find that position learning networks rely more on local computations and form circuits associated with particular ‘tasks’ after training. Not only do our findings have implications for our understanding of the brain but also, suggest new ways of studying neural networks and potential directions for future neuromorphic implementations.
2 Results
2.1 Position learning links structure to task
We first introduce a novel neuron position learning algorithm, in which neurons move in space so as to minimise the objective function. This makes use of the fact that delays between neurons can augment performance. Intuitively, a neuron with a great spatial separation should therefore correspond to a longer delay. Recently, various delay learning algorithms have been derived, and shown to be an effective way of improving the performance of SNNs (Hammouamri et al., ; Sun et al., 2023; Göltz et al., 2024; Mészáros et al., 2025). Mészáros et al. (2025) was the first to introduce delay learning for recurrent connections. Spatial embeddings become particularly interesting with bidirectional connectivities, thus we derived our equations in a recurrent framework. In a 2-dimensional space, we define the delay between neuron $j$ and $i$ as:
$d_{ji}=\sqrt{(x_{i}-x_{j})^{2}+(y_{i}-y_{j})^{2}},$ (1)
where $(x_{i},y_{i})$ and $(x_{j},y_{j})$ denote the positions of neurons $i$ and $j$ in a 2-dimensional space. Thus, the spike arrival time at neuron $m$ becomes:
$t_{lm}\equiv t_{l}+d_{mn(l)}=\\ t_{l}+\sqrt{(x_{m}-x_{n(l)})^{2}+(y_{m}-y_{n(l)})^{2}},$ (2)
where $n(l)$ is the neuron, that fired the $l$-th spike, at time $t_{l}$. Similarly to the original algorithm (Wunderlich & Pehle, 2021) and its extensions (Mészáros et al., 2025), we also work with Leaky Integrate-and-Fire (LIF) neurons, where each neuron has an input current $I$, and a membrane voltage $V$.
[…]
$\frac{d\mathcal{L}}{dx_{i}}=\sum_{j}\left(\frac{d\mathcal{L}}{dd_{ji}}\frac{dd_{ji}}{dx_{i}}+\frac{d\mathcal{L}}{dd_{ij}}\frac{dd_{ij}}{dx_{i}}\right)$ (3)
From (Mészáros et al., 2025), equation (6), we know
$\frac{d\mathcal{L}}{dd_{ji}}=-w_{ji}\sum_{t_{k}\in\,\parbox{18.49411pt}{\scriptsize spikes from\ $i$}}(\lambda_{I}-\lambda_{V})_{j}\Big|_{t_{k}+d_{ji}}.$ (4)
and analaguosly for $\frac{d\mathcal{L}}{dd_{ij}}$ and by taking the derivative of equation (1), we get
$\frac{dd_{ji}}{dx_{i}}=\frac{x_{i}-x_{j}}{d_{ji}}=\frac{dd_{ij}}{dx_{i}}.$ (5)
Taken together,
[…]
Here, $\mathcal{L}$ is the loss and $w_{ij}$ is the synaptic weight. $\lambda_{I}$ and $\lambda_{V}$ denote the backwards sensitivities corresponding to the input current and membrane voltage, respectively. We observe that these updates are a sum of of all presynaptic and postsynaptic delay gradients, normalised by $(x_{i}-x_{j})/d_{ji}$. We thus arrive at a learning algorithm with the same computational benefits as before – neurons have the same computational requirements for the forward and backward passes, the costly synaptic updates only happen at spike times in both directions.
Temporally complex tasks benefit from introducing delays (Hammouamri et al., ; Habashy et al., 2024; Sun et al., 2023; Göltz et al., 2024) and adding delays to recurrent connections is particularly beneficial for small networks (Mészáros et al., 2025; Queant et al., 2025). This raises the question: do spatial embeddings serve as a useful inductive bias by defining delays as the Euclidean distance between two neurons? If so, what would otherwise be considered a constraint could be leveraged.
We investigate this by training an SNN with learnable positions on the SHD classification task (Cramer et al., 2020). By learning positions instead of synaptic delays, we reduce the number of free parameters from $n^{2}$ to $n\times d$, where $n$ is the number of neurons, and $d$ is the number of dimensions. Furthermore, this set-up yields delay distributions which inherit the following properties from the Euclidean distance:
  • Zero diagonals: Autapses take effect in the next timestep
  • Symmetry: Delay from neuron $i$ to neuron $j$ is the same as from $j$ to $i$
  • Triangle inequality: The ‘quickest’ path to another neuron is always the direct path
To ascertain that position learning is useful, we first compare position learning networks of various sizes and embeddings of various dimensions, against networks with non-spatially constrained synaptic, and axonal delays, and networks without any delays. We could have additionally compared against networks with fixed random delays, but it raises the question of how to initialise these delays. In recurrent connections, this is a particularly complex and understudied question. We illustrate the comparisons of the networks in the Appendix.
With a fixed number of trainable parameters, we observe that networks with delays always outperform networks without them. As network size (i.e. number of hidden neurons) increases, the performance gaps decrease. This is not surprising, as all architectures have recurrent connections, which, similar to delays, can serve as a temporal memory. As the network gets more and more neurons to store information in, delays become less and less necessary. Spatially embedded networks perform similarly to networks with unconstrained delay learning. Axonal delays, however, perform at least as well as the synaptic delays derived from the Euclidean distances. Note, that the error bars are large because we study small networks, to make sure that there is a clear benefit of delays.
The formulation does not allow for neurons to be placed on top of each other, but in all cases we initialise the shape such that distances are close to 0 (and for non-spatial architectures we initialise all delays to 0).
We conclude that recurrent delay learning for positions improves task performance – a necessity for the rest of our studies. From here, we focus on one network size and embedding dimension – 2D networks with a hidden layer size of 128 – since we observed no fundamental differences between network sizes and embedding dimensions.
Similarly to Achterberg et al. (2023) we also study the effects of L1 regularisation, and L1 regularisation scaled with neuron distance. For our studies, we force neurons into an n-dimensional sphere (so that the effect of the length-scaled regularisation will be constrained), with a diameter of 62, such that the maximum possible delay is equal to the maximum delay allowed on Loihi 2 (Davies et al., 2021). Since regularisation will not drive weights to become exactly 0, we tune our hyperparameters based on validation accuracies obtained after dropping the 90% lowest magnitude connections.
2.2 Nearby neurons become functionally connected
We looked at the relationship between neuron positions in the hidden layer and the inputs they preferentially respond to, post-training. First, we tried to predict the X, Y positions of hidden neurons from their corresponding 700-dimensional input to hidden weight vectors using ridge regression. We found that – based on the observed R2 scores shown in Table 1 – we can do this quite accurately, suggesting that neurons with similar input patterns are physically close as well.
Init No reg. L1 L1+dist.
X $0.07$ $0.82\pm 0.1$ $0.72\pm 0.1$ $0.6\pm 0.1$
Y $0.07$ $0.78\pm 0.1$ $0.59\pm 0.1$ $0.61\pm 0.1$
Table 1: R2 scores
We also looked at the centre of mass for each input neuron, by taking the mean of all neuron positions weighted with the corresponding synaptic strength, and, as figure 1 shows, neurons in both networks clearly separate depending on which ‘part’ of the input they focus on (i.e higher or lower frequencies).
Figure 1: A Input example from the SHD dataset. B Example of preferred positions of input neurons based on the learnt synaptic strengths. C Example of preferred position of input neurons, based on their activity. D Examlpe of preferred position over time. The depicted plots are from networks trained with L1+distance cost regularisation, but we observe such spatial patterns under all training circumstances.
Next, we measured the average (over input samples) spiking activity of each hidden neuron in a 5 ms window after each input spike and then took the average of this 3-dimensional tensor (input neuron $\times$ hidden neuron $\times$ window) over time. Note that this measure is different from measuring firing rates, as postsynaptic spikes need to happen in a small window after presynaptic spikes. Doing this, we again end up with a matrix that is the same shape as the input-to-hidden weight matrix. We can now again take the weighted average of neuron position according to these values. For results, see Figure 1. Instead of taking the average over time, we can also measure the average sensitivity for each timestep and look at where computation is happening for each input neuron. We observe that over time, computation ”moves” towards the centre and spreads out again by the end of the trial.
2.3 Higher modularity and lower entropy through distance cost
Next we studied various metrics in trained networks. Since the weights are regularised, we can expect many weights to be close to but not exactly zero. Thus, we measure the weighted modularity metric (for definition see Appendix A). In Figure 2, we can observe that the peak modularity Q is achieved relatively early on in training, and the distance cost needs to be introduced, for higher modularity (i.e position learning will not inherently lead to more modular weight matrices).
Shannon entropy (for definition see Appendix A) quantifies the amount of unpredictability. A highly predictable system has low entropy and will exhibit a degree of clustering around particular weight values. An unpredictable system has high entropy and will be more uniform in its distribution. While the previously introduced modularity metric Q shows the extent to which the network can be partitioned into subcommunities, it does not measure the degree of uncertainty of the weight distribution. In Figure 2 we observe, that position learning networks achieve lower entropy than non-spatial learning networks, but introducing the distance cost decreases entropy even further.
In this work, we are interested in the effect of space on network structure, and thus only introduced the distance cost to the regularisation term. Achterberg et al. (2023) and Sheeran et al. (2024) also introduced a communicability cost, which we did not employ here. However, we can still measure the communicability entropy, similarly to Sheeran et al. (2024) (for definition see Appendix A). In figure 2, we can observe that networks with the introduced distance cost yield a lower entropy. Given that weight and communicability entropy are both lower for positions learning networks, we can conclude that these architectures yield a small number of highly communicable connections, with an increasingly ordered topology.
We also measured the small-worldness on the connectivity matrix with the 90% lowest magnitude connections dropped. Again, we observe that the distance cost term increases the metric. Interestingly, without the distance cost, we observe lower small-worldness in position learning networks, compared to synaptic delay learning.
We can also visualise the learnt network shapes together with the corresponding synaptic strengths and figure 2 compares these with and without the spatial cost. We can observe that the network shapes end up looking similar (with neurons moving close to the perimeter), but without the spatial cost, we end up having stronger long-range connections. There is no direct driving force on the neuron positions, and thus they try to maximise the possible delays, but through the distance-based regularisation, the connectivity patterns are more local.
Figure 2: Network shapes without (A) and with spatial cost (B). Line thickness corresponds to synaptic strength. Neurons move close to the perimeter in both cases, but when introducing the distance cost, we have fewer long strong connections. C-F Training dynamics over epochs. D shows that the network with the introduced cost term achieves a higher modularity than the other two. In E we observe that spatial networks achieve a lower Shannon entropy in their weights, regardless of the cost term. On C we see that if calculate the Shannon entropy on the communicability matrix derived from the weight matrix, the network with the distance cost has a lower value. On C we observe that introducing the cost term yields higher small-worldness.
In this section, we have applied the metrics used in Sheeran et al. (2024) to our models. We observe the same distinctions between the non-spatial and spatial architectures, but to a smaller extent than what they showed. While it is possible that this is due to the introduction of position learning and delays, it is more likely that it is thanks to the fact that we tuned our hyperparameters for high accuracy (after pruning), and weaker regularisation. Furthermore, comparing networks with and without a distance cost with the same regularisation strength is not a fair setup. As neurons can change positions, this could too heavily influence the network dynamics, as the regularisation term is continuously changing not only according to the weights, but according to the network shape as well. Additionally, Sheeran et al. (2024) and Achterberg et al. (2023) added a communicability matrix to the regularisation term, which we do not study here.
2.4 Unconstrained position learning converges on more local solutions
Figure 3: A pruning effects on classification accuracy on SHD. Pruning is less harmful for the position learning network. B delay distributions post training. For the non-spatial network, most delays are still zero after training, but we also end up with a few very long delays that are non visible since they are significantly outnumbered by the zero delays. For the spatial network delays become spread out in the lower range. C delay length based pruning effects on modularity. Position learning network achieves higher modularity. D delay length-based pruning effects on small-worldness. Position learning network achieves higher small-worldness.
We also study networks with no regularisation, which, of course, is not biologically realistic, as connecting to nearby neurons in the brain is always ‘cheaper’. We were interested, however, in what solutions the network finds under no constraints, and if we still converge on solutions where certain synapses are more important than others. We found that ‘classic’ weight magnitude-based pruning and random pruning had very similar effects on both non-spatially embedded networks and those where the positions were learnt. Taking inspiration from biology and emphasising short connections, we pruned away the top $x\%$ (ranging from $5\%$ to $30\%$) longest connections. The learnt delay distributions of non-spatial and spatial networks seem to also hint (see Figures 3), that our pruning studies show that spatial networks converge on solutions that are more reliant on short local connections (see Figure 3). Given that this was not enforced through any explicit regularisation, this is a surprising finding, particularly since long delays tend to be beneficial for temporal tasks (Sun et al., 2025). These findings not only align with neuroscience properties but they have implementation benefits as well. In digital neuromorphic systems, delays are usually implemented through costly dynamic memory buffers, so shortening delays can be beneficial. Furthermore, most neuromorphic chips also have an upper bound on delays (Davies et al., 2021). Given that random and magnitude-based pruning did not show such differences, seemingly spatial networks are not simply more robust but are more ‘local’. We also tested robustness through neuron ablation studies, and adding noise to delays as well and we did not observe a significant difference between spatial and non-spatial architectures.
Since delay length based pruning has less impact on the performance of spatial architectures, we were curious to see how the network topology changes after such pruning. We again study modularity in our networks, and also small-worldness. Note, that neither of them take physical distances between nodes into account. Figure  3 shows the effect of delay length pruning on modularity and small-worldness. As a baseline, we calculated the modularity of a random matrix, and found that spatial architectures show a clear increase in modularity with pruning, while non-spatial architectures show a similar trend as pruning a random matrix. Similarly, small-worldness increases more rapidly in spatial networks (with the same connectivity ratio). We observe the higher accuracy, modularity and small-worldness after pruning in spatial architectures, without explicitly enforcing any of these structures to emerge.
3 Discussion
Here we introduced a method to study the intertwinement of space and time in neural networks. While neuron position learning has been explored before, prior studies have mainly emphasized parameter efficiency (Erb et al., 2025) or performance improvements (Vasilache et al., 2025) on machine learning benchmarks. Here, we instead focus on how the structure or learnt networks changes, when space and time are intrinsically connected.
We find that neuron positions in the hidden layer correlate with input-layer connectivity patterns. However, in our simple classification setting, it remains unclear what functional roles clusters of nearby neurons serve. Long delays between modules specialized for different subtasks may be advantageous when temporal alignment across these modules is required for solving a more complex task. Extending this analysis to modular tasks (Béna & Goodman, 2025) with an added temporal structure, to multimodal settings, or multitask problems (Vafidis et al., 2024) presents an interesting avenue for future work.
In our examples we focus on the inputs’ representation in the hidden layer. Of course it would be interesting to see if neurons (or even groups of neurons) specialise on particular inputs, but since we are compressing from the input to the hidden layer (700 input neurons to 128 neurons), we would not expect specialisation like that to emerge. As mentioned before, we study small networks, to make sure that delays become useful for for the task. We could also of course constrain the networks in other ways such as by constraining neurons to only spike once. Another solution would be to turn to tasks where we have a small input population with complex information, such that recurrent delays are still crucial even when we expand from the input to the hidden layer.
Achterberg et al. (2023) investigated networks with fixed neuron positions. If positions were instead allowed to be optimized, all neurons would likely converge toward the centre. In such a case, the benefits of modularity and small-world structure would vanish, since distance acts only as a cost. Once proximity carries no penalty, the network is free to adopt any connectivity pattern needed to solve the task. In contrast, by introducing a benefit of distance, our framework prevents this collapse and encourages spatially distributed structures. Similarly, in biological neural networks, most connections are short-range, yet the overall average connection length remains higher than the physical minimum (Bassett et al., 2010). During development, neuronal migration (de Graaf-Peters & Hadders-Algra, 2006) contributes to functional circuit formation. However, biological constraints such as each neuron occupying a physical volume (volume exclusion) prevents neurons from being too close to one another. In our model, neurons can theoretically be arbitrarily close, provided they are not coincident. Introducing an explicit repulsive force between two neurons would be straightforward to formalize but challenging to extend to more complex circuits—offering another interesting direction for future work. Furthermore, now we even have studies that track neurons position changes over time (Akarca et al., 2025), allowing for direct comparisons between the presented algorithm and empirical data.
As we have shown, with position learning, networks tend to converge to more local solutions, where the core computations occur within small-world and modular topologies. While these findings are interesting from a neuroscience perspective, such solutions may also be beneficial for neuromorphic implementations. For example, the chip introduced by Dalgaty et al. (2024) relies on small-world connectivity. Weber et al. (2025) demonstrated that the structural plasticity algorithm DEEP R (Bellec et al., 2017) can be extended to account for system constraints (e.g., small-world topology). Inspired by this, introducing distance awareness with position learning to DEEP R could yield interesting network geometries, as neuron positions would no longer need to be constrained to a hypersphere (as we did here) to comply with system constraints.
Keller et al. (2024) outlines the importance of spacetime neural representation both for theoretical neuroscience and improved generalisation and long-term working memory in AI. Travelling waves have been shown to facilitate global information integration (Jacobs et al., 2025). Izhikevich & Hoppensteadt (2009) even proposed to implement such methods without synapses (or in other other words, fully connected networks with uniform synapses), showing that such set up can perform temporal pattern recognition, reverberating memory, temporal signal analysis and basic logical operations. Because distance-dependent delays in large spiking networks naturally give rise to such waves (Davis et al., 2021), investigating position learning in this context is a particularly promising direction.
Appendix A Methods
For this work, the original EventProp backward dynamics (Wunderlich & Pehle, 2021; Nowotny et al., 2025) do not change. The weight update rule with delays (Mészáros et al., 2025) is defined as
$\frac{{\rm d}\mathcal{L}}{{\rm d}w_{ji}}=-\tau_{s}\sum_{t_{k}fromi}\lambda_{I,j}\Big|_{t_{k}+d_{ji}},$ (7)
where $\mathcal{L}$ is the loss, $w_{ji}$ is the synaptic weight from neuron $i$ to $j$, $\tau_{s}$ is the synaptic time constant, $t_{k}$ is the k-th spike, $\lambda_{I}$ is the input current backward dynamic for neuron $j$, and $d_{ji}$ is the delay between neuron $j$ and $i$. If we apply an L1 regularisation term, the update rule becomes
$\frac{{\rm d}\mathcal{L}}{{\rm d}w_{ji}}=-\tau_{s}\sum_{t_{k}fromi}\lambda_{I,j}\Big|_{t_{k}+d_{ji}}+\Lambda_{1}sign(w_{ji}),$ (8)
where $\Lambda_{1}$ is the regularisation strength. If we introduce the distance cost as well, the equation is
$\frac{{\rm d}\mathcal{L}}{{\rm d}w_{ji}}=-\tau_{s}\sum_{t_{k}fromi}\lambda_{I,j}\Big|_{t_{k}+d_{ji}}+\Lambda_{1}sign(w_{ji})d_{ji}.$ (9)
We define the modularity statistic Q as
$Q=\frac{1}{l}\sum_{i,j\in N}\left(a_{i,j}-\frac{k_{i}k_{j}}{l}\right)\delta_{m_{i}m_{j}},$ (10)
where $l$ is the total number of synapses (or the the sum of all synapse strengths in the weighted case), $N$ is the total number of neurons, $a_{ij}$ is the connection status between neuron $i$ and $j$ (or the synapse strength in the weighted case), $k_{i}$ is the number of synapses of neuron $i$, and $m_{i}$ is the module containing neuron $i$.
We define small-worldness as
$\sigma=\frac{c/c_{rand}}{l/l_{rand}},$ (11)
where $c$ and $c_{rand}$ are clustering coefficients, and $l$ and $l_{rand}$ are the characteristic path lengths of the tested network and the mean of $1000$ random networks with the same size and density. In the fully connected network $\sigma\approx 1$.
The Shannon entropy of the weight matrix $W$ is defined as
$H(W)=-\frac{1}{N}\sum_{i=1}^{N}\frac{w_{ij}}{\sum_{i=1}^{N}w_{ij}}log_{2}\Big(\frac{w_{ij}}{\sum_{i=1}^{N}w_{ij}}\Big),$ (12)
We define the communicability matrix as
$C=e^{-\frac{1}{2}}We^{-\frac{1}{2}},$ (13)
and we can calculate the communicability entropy with the combination of the previous two equations,
$H(W)=-\frac{1}{N}\sum_{i=1}^{N}\frac{C_{ij}}{\sum_{i=1}^{N}C_{ij}}log_{2}\Big(\frac{C_{ij}}{\sum_{i=1}^{N}C_{ij}}\Big).$ (14)
For this work we used the mlGeNN library (Knight & Nowotny, 2023). We implemented support for weight regularisation, the position gradient update function, and axonal delays.
Appendix B Architecture comparisons
Figure 4: Architecture comparisons